Consider the system of equations
3x+2y=17
x-y=-1

Answer by xcentaur(357)

3x+2y=17.....[1]
x-y=-1.......[2]


multiply eqn[2] by 2
3x+2y=17
2x-2y=-2


Add up both the equations:
3x+2y+(2x-2y)=17+(-2)
5x=17-2
5x=15
x=3


x-y=-1
-y=-1-x
-y=-(1+x)
y=1+x
y=1+3
y=4


x=3,y=4


-xC